### v1

# uGMM-NN: Univariate Gaussian Mixture Model Neural Network

Abstract
This paper introduces the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel neural architecture that embeds probabilistic reasoning directly into the computational units of deep networks. Unlike traditional neurons, which apply weighted sums followed by fixed nonlinearities, each uGMM-NN node parameterizes its activations as a univariate Gaussian mixture, with learnable means, variances, and mixing coefficients. This design enables richer representations by capturing multimodality and uncertainty at the level of individual neurons, while retaining the scalability of standard feedforward networks. We demonstrate that uGMM-NN can achieve competitive discriminative performance compared to conventional multilayer perceptrons, while additionally offering a probabilistic interpretation of activations. The proposed framework provides a foundation for integrating uncertainty-aware components into modern neural architectures, opening new directions for both discriminative and generative modeling.
Keywords: probabilistic neurons, deep learning, probabilistic circuits, Gaussian mixture, uGMM, uGMM-NN
1 Introduction
Deep neural networks have transformed machine learning, excelling in tasks such as image classification and natural language processing through hierarchical feature learning [1]. However, traditional neurons, which compute deterministic weighted sums followed by nonlinear activations (e.g., ReLU, sigmoid), struggle to model uncertainty or multimodal distributions prevalent in real-world data. This limitation has historically been addressed by probabilistic graphical models, such as Bayesian Networks [2] and Markov Random Fields [3], which offer robust frameworks for uncertainty quantification and complex dependency modeling [4]. These models provide a strong conceptual foundation, but often lack the deep hierarchical feature learning capabilities of modern neural networks.
A key research focus has therefore been on bridging the gap between these two paradigms. This has led to approaches that incorporate the probabilistic principles of graphical models directly into deep learning architectures. For example, Bayesian Neural Networks (BNNs) embed uncertainty into network weights [5], while Probabilistic Circuits (PCs), including Sum-Product Networks (SPNs) [7, 8, 9], are deep probabilistic models that build on a formal probabilistic structure, fusing the representational power of graphical models with the hierarchical feature learning of neural networks.
In contrast, this paper introduces a novel approach by embedding a univariate Gaussian Mixture Model (uGMM) directly into the network’s computational units, enabling each neuron to represent a multimodal univariate Gaussian distribution, thereby directly addressing the limitations of deterministic neurons from a new perspective. While our primary focus is on discriminative applications where the goal is to predict a class label, the probabilistic backbone of our model shares conceptual similarities with probabilistic circuits. This suggests that when trained generatively, the model could potentially learn complex joint distributions. We evaluate uGMM-NN on two representative datasets. On the image dataset MNIST, we compare its discriminative performance against a standard Multi-Layer Perceptron (MLP). On the tabular dataset, Iris, we highlight the model’s ability to capture joint distributions for inference, drawing conceptual connections to probabilistic graphical models and probabilistic circuits.
2 Related Work
uGMM-NN relates to several paradigms in probabilistic graphical modelling and deep probabilistic (differentiable) learning. Bayesian Neural Networks (BNNs) model uncertainty over weights [6], providing a principled route to calibrated predictions, though often at the cost of heavier inference. Probabilistic Circuits, such as Sum-Product Networks [7], enable tractable inference for complex distributions [8, 9]. These models provide a framework for explicitly representing probability distributions with guarantees on tractability, a concept uGMM-NN draws inspiration from by embedding a tractable probabilistic model within each neuron.
In parallel, Kolmogorov–Arnold Networks (KAN) [10] pursue interpretability and flexibility by replacing the linear affine maps of MLPs with learnable univariate functions along edges. KAN thus substitutes the standard transformation:
$x\mapsto Wx+b$
with a composition of spline-like functions, yielding more transparent functional dependencies than opaque weight matrices. uGMM-NN makes an analogous replacement at the unit level: instead of a linear neuron followed by a pointwise nonlinearity, each computational unit is a univariate Gaussian Mixture Model (uGMM) that outputs responsibilities and mixture statistics. uGMM-NN swaps linear neurons for probabilistic neurons, yielding interpretable, multimodal, and uncertainty-aware intermediate representations.
This substitution brings two benefits central to our aims. First, interpretability: mixture components act as localized, semantically meaningful “submodes,” exposing how a unit partitions its input space, akin to KAN’s readable univariate maps, but now in explicitly probabilistic terms. Second, expressivity under uncertainty: mixtures capture multimodality and uncertainty within each unit, rather than deferring uncertainty to the weight space (as in BNNs) or to global circuit structure (as in SPNs and probabilistic circuits). In short, like KAN replaces linear MLP layers with functional mappings, uGMM-NN replaces them with probabilistic mixture units, yielding a network whose intermediate computations are distributions, interpretable, multimodal, and uncertainty-aware rather than single deterministic activations.
3 uGMM-NN
The uGMM-NN replaces standard MLP neurons with probabilistic units, each modeling a univariate Gaussian mixture over its inputs. This design enables the network to capture multimodal behavior and input uncertainty at the level of individual neurons, while remaining fully differentiable and trainable with standard supervised losses. Conceptually, each uGMM neuron serves as a direct replacement for a traditional linear neuron followed by a pointwise nonlinearity, providing richer intermediate representations.
3.1 Univariate GMM Nodes
In traditional multilayer perceptrons (MLPs), each neuron computes a weighted sum of inputs followed by a nonlinear activation function such as ReLU or sigmoid. These scalar activations are then passed forward to the next layer. In contrast, each neuron in uGMM-NN computes the log-density of a univariate Gaussian Mixture Model (uGMM), making the neuron’s output a probabilistic quantity rather than a deterministic scalar.
Let $\mathbf{x}=[x_{1},\dots,x_{N}]$ denote the inputs to neuron $j$ from the previous layer, where $N$ is the number of neurons in that layer. In uGMM-NN, each neuron $j$ defines a univariate Gaussian mixture over a latent variable $y$, parameterized by learnable means $\mu_{j,k}$, variances $\sigma^{2}_{j,k}$, and mixing coefficients $\pi_{j,k}$ for each input $x_{k}$:
$P_{j}(y)=\sum_{k=1}^{N}\pi_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma^{2}_{j,k}),\quad\text{with}\quad\sum_{k=1}^{N}\pi_{j,k}=1.$
Here, each input contributes a component to the mixture distribution. The number of components $N$ is thus determined by the width of the previous layer, not treated as an independent hyperparameter.
Expanding the Gaussian explicitly, the probability density becomes:
$P_{j}(y)=\sum_{k=1}^{N}\pi_{j,k}\cdot\frac{1}{\sqrt{2\pi\sigma^{2}_{j,k}}}\exp\left(-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right).$
The neuron’s output is then defined as the log-density:
$\log P_{j}(y)=\log\left(\sum_{k=1}^{N}\exp\left[\log\pi_{j,k}-\frac{1}{2}\log(2\pi\sigma^{2}_{j,k})-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right]\right),$
which is computed using the log-sum-exp trick for numerical stability.
We refer to this log-density output $\log P_{j}(y)$ as the neuron’s activation [^1] In standard neural networks, an “activation” refers to a scalar value resulting from a nonlinear transformation of a weighted sum. In uGMM-NN, each neuron instead outputs the log-density of a mixture distribution, which we treat as the probabilistic analogue of a traditional activation. These log-densities are passed to the next layer, participate in gradient flow, and play the same architectural role.. This probabilistic activation is propagated forward through the network and forms the basis for both classification and likelihood-based training. At subsequent layers, the inputs to each uGMM neuron are the log-density activations produced by the preceding layer, ensuring that a consistent probabilistic representation is maintained throughout the network.
[…]
Figure 1: Schematic of a uGMM-NN neuron. Each neuron models a uGMM with NN components based on inputs from the previous layer, with each connection parameterized by mean ($\mu_{j,k}$), standard deviation ($\sigma_{j,k}$), and mixing coefficient ($\pi_{j,k}$), outputting the log probability $\log P_{j}(y)$.
3.2 Probabilistic Interpretability of Neurons
Unlike standard MLP neurons, which produce deterministic scalar activations, uGMM-NN neurons output the log-density of a univariate Gaussian mixture. This probabilistic quantity not only propagates forward through the network but also encodes interpretable structure at the unit level.
Figure 2 illustrates a single uGMM neuron with three mixture components. Each component is parameterized by a mean $\mu_{j,k}$, variance $\sigma_{j,k}^{2}$, and mixing coefficient $\pi_{j,k}$. Together, these define the neuron’s overall probability density $P_{j}(y)$, with the dashed black line representing the full mixture. The neuron’s output is the log-density $\log P_{j}(y)$, which we interpret as its activation.
This formulation enables an intuitive understanding of what the neuron ”detects.” For example, a component with a high mixing coefficient $\pi_{j,k}$ indicates that the corresponding input $x_{k}$ plays a dominant role in shaping the neuron’s density. Similarly, the spread of the Gaussian (via $\sigma_{j,k}^{2}$) reflects how confidently the neuron attributes input mass to a given region of its input space. From this perspective, each neuron defines a probabilistic partition over its input domain, where mixture components correspond to interpretable submodes of the neuron’s response. This is in contrast to traditional activations, which typically lack such localized semantics. As a result, uGMM-NN offers built-in interpretability, where the internal representation of each layer is not just a scalar value but a structured probability landscape.
Figure 1 (introduced in the previous section) situates this mechanism within the full network, illustrating how log-densities are propagated forward through successive layers. Together, Figures 1 and 2 demonstrate how probabilistic activations are formed and refined across the network. This property is particularly advantageous in applications such as healthcare or autonomous systems, where interpretability and uncertainty quantification are essential.
[…]
Figure 2: Illustration of a single uGMM neuron with three univariate Gaussian mixture components. The means ($\mu_{j,k}$), variances ($\sigma_{j,k}^{2}$), and mixing coefficients ($\pi_{j,k}$) define the neuron’s probability density $P_{j}(y)$. The dashed black line shows the combined mixture density.
3.3 Network Structure and Training
The uGMM-NN is organized in a standard feedforward manner with input, hidden, and output layers. Each neuron in a hidden layer computes a univariate Gaussian mixture over the activations of the previous layer, producing a log-density that serves as input for the next layer. Importantly, the number of mixture components in each neuron is automatically determined by the width of the preceding layer, ensuring a direct mapping from inputs to probabilistic components.
Stacking multiple uGMM layers allows the network to build hierarchical probabilistic representations, capturing increasingly complex and multimodal patterns in the data. Unlike traditional MLPs, where each neuron performs a 1D affine transformation followed by a nonlinearity and outputs a deterministic scalar, uGMM neurons operate on univariate inputs and output log-probability values under a learned Gaussian mixture. These can be viewed as 1D probabilistic operations, analogous in shape and dimension to their MLP counterparts. By focusing on univariate GMMs at each neuron, uGMM-NN avoids the exponential complexity often associated with multivariate mixture models, allowing it to scale efficiently to deep architectures. This architectural simplicity enables practical training and inference on large datasets while retaining the expressive power of probabilistic modeling.
For discriminative tasks, the output (root) layer log-densities can be converted into class probabilities using the softmax function, after which training minimizes the standard cross-entropy loss.
Gradients are backpropagated through the log-density computations using automatic differentiation, and training is performed with gradient-based optimizers such as Adam or SGD. To improve numerical stability, the log-sum-exp trick is employed when computing gradients through the mixture sums, preventing overflow or underflow in deep networks.
Despite the added complexity of per-neuron mixture computations, uGMM layers are naturally amenable to GPU parallelization. Each neuron’s responsibilities and log-density contributions are computed using vectorized tensor operations over batches, neurons, and mixture components. Specifically, the core computation, which evaluates the log-density of each input under a learnable mixture of Gaussians, is implemented as a fully batched operation using PyTorch, with dropout applied directly to the mixture components during training. The use of logsumexp ensures numerical stability, while broadcasting enables efficient computation across batch dimensions. As a result, uGMM-NN achieves training and inference speeds comparable to standard multilayer perceptrons despite the probabilistic overhead [12, 13].
3.4 Dropout in uGMM-NN
To improve generalization and prevent overfitting, dropout can be applied to uGMM neurons in a manner analogous to standard MLP layers. Specifically, a Bernoulli mask is applied independently to each mixture component $k$ of neuron $j$:
$d_{j,k}\sim\text{Bernoulli}(1-p),$
where $p$ is the dropout probability. Components that are ”dropped” are effectively removed from the mixture by setting their log-density contribution to -$\infty$. The resulting log-density with dropout is
$\log P_{j}(y)_{\text{dropout}}=\log\sum_{k=1}^{N}\pi_{j,k}\,d_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma_{j,k}^{2}),$
where $\pi_{j,k}$, $\mu_{j,k}$, and $\sigma_{j,k}^{2}$ are the mixture coefficients, means, and variances of the $k$-th component of neuron $j$, respectively. Dropout is applied only during training and is disabled during inference, allowing the full mixture to contribute to the neuron’s output.
3.5 Universal Approximation Perspective
Each uGMM neuron outputs a log-density defined by a mixture of univariate Gaussians. Since a mixture of Gaussians can approximate any smooth univariate probability density on a compact domain to arbitrary accuracy, uGMM neurons are, in principle, universal approximators of univariate densities.
When stacked in a feedforward network, the resulting uGMM-NN can approximate complex multivariate distributions over the inputs. For discriminative tasks, this implies that the network can represent any conditional probability distribution $P(y\mid x)$ given sufficient width (number of neurons per layer) and depth (number of layers).
In comparison, standard MLPs and Kolmogorov–Arnold Networks (KANs) provide universal approximation guarantees for deterministic functions. uGMM-NN extends this notion to probabilistic, multimodal representations, combining expressive power with interpretable uncertainty through mixture parameters $(\pi_{j,k},\mu_{j,k},\sigma_{j,k}^{2})$.
The uGMM-NN architecture and all experiments were implemented in PyTorch. For full reproducibility, the code corresponding to this work is available via a Zenodo release (initially published on July 20, 2025) [12], and the latest development version can be accessed on GitHub [13].
4 Experiments
We empirically evaluate the uGMM-NN on classification benchmarks, comparing it against a standard Feedforward Neural Network (FFNN) with equivalent architecture. The FFNN consists of linear layers with ReLU activations, while the uGMM-NN replaces each neuron with a uGMM node as described in Section 3.1. To ensure a fair comparison, both models were matched in hidden layer sizes and total depth. Dropout was applied in all models to reduce overfitting, and hyperparameters were selected based on validation performance.
4.1 Datasets
We considered two widely used benchmarks:
  • MNIST [14]: A dataset of 70,000 grayscale images (60,000 train, 10,000 test) of handwritten digits ($\times28$ pixels), with 10 output classes.
  • Iris [15]: A classic dataset of 150 samples with 4 numerical features and 3 output classes, corresponding to different iris species.
4.2 Training Setup
All models were trained using the Adam optimizer with default parameters ($\beta_{1}=0.9,\beta_{2}=0.999$), and an initial learning rate tuned in $\{10^{-4},10^{-3},10^{-2}\}$. Batch normalization was applied in the FFNN baselines where applicable.
Both models were trained for $100$ epochs with the following learning rate schedule:
$\texttt{MultiStepLR}(\texttt{milestones}=[20,45,60],\ \gamma=0.1).$ (1)
The FFNN used an initial learning rate of $10^{-3}$ and applied dropout with probability $p=0.5$ on the 64-unit layer. The uGMM-NN used a larger initial learning rate of $10^{-2}$, reflecting the faster adaptation required by mixture parameters, and applied dropout with probability $p=0.3$ on the 128-unit layer.
For uGMM-NN, the mixture parameters $(\mu_{j,k},\sigma_{j,k}^{2},\pi_{j,k})$ were initialized randomly, and gradients were backpropagated through the log-density computations using automatic differentiation. On the Iris dataset, uGMM-NN was trained generatively to maximize the joint likelihood of features and labels, with predictions obtained by posterior inference. On MNIST, both uGMM-NN and FFNN were trained discriminatively using cross-entropy loss at the output layer.
4.3 Performance
Results are reported in Table 1.
On the Iris dataset, uGMM-NN was trained generatively to maximize the joint likelihood of features and labels. Predictions were obtained via posterior inference:
$\hat{y}=\arg\max_{c}P(y=c\mid\mathbf{x})=\arg\max_{c}P(y=c,\mathbf{x}),$
where $\mathbf{x}$ is the feature vector. The FFNN baseline was trained discriminatively with cross-entropy loss. Both models achieved perfect classification accuracy, illustrating that uGMM-NN can perform competitively on tabular datasets even when trained generatively. While the Iris dataset is relatively simple, these results motivate future evaluation of uGMM-NN on more complex tabular benchmarks.
On MNIST, both models converged to strong performance after 100 epochs. The FFNN achieved a stable test accuracy of $98.21\%$ (error rate $1.79\%$) in the final training window (epochs 90–100). The uGMM-NN attained a slightly lower but competitive test accuracy of $97.74\%$ (error rate $2.26\%$) across the same interval. This indicates that while the FFNN maintains a small advantage in raw discriminative accuracy, the uGMM-NN remains within $0.5\%$ of the baseline while offering additional benefits such as uncertainty quantification and a probabilistic interpretation of hidden activations.
Table 1: Test accuracy and training loss for uGMM-NN vs. FFNN across benchmark datasets.
Dataset Model Test Accuracy (%) Training Loss
Iris FFNN 100.0 $\pm$ 0.0 Cross-Entropy
Iris uGMM-NN 100.0 $\pm$ 0.0 NLL (Generative)
MNIST FFNN 98.21 Cross-Entropy
MNIST uGMM-NN 97.74 Cross-Entropy
5 Discussion and Limitations
The uGMM-NN integrates probabilistic modeling into deep learning, offering interpretable uncertainty quantification through its mixture components. Compared to standard MLP neurons, each uGMM neuron introduces additional parameters, specifically, means, variances, and mixing coefficients for each input, enabling richer, multimodal representations at the cost of increased model size.
Tying the means $(\mu_{j,k})$ of each Gaussian component to the corresponding input activation $(x_{k})$, i.e., setting $\mu_{j,k}=x_{k}$, is a straightforward way to reduce the number of learnable parameters. In this case, the neuron only needs to learn the mixing coefficients $(\pi_{j,k})$ and variances $(\sigma_{j,k}^{2})$, simplifying the model while reducing expressivity, since the neuron can no longer shift submodes independently of the inputs.
A remaining limitation arises when using the model for generative inference, in a manner similar to Bayesian networks, or probabilistic circuits [8, 9]. Computing the Most Probable Explanation (MPE) for a full generative network is non-trivial, and currently no efficient algorithm exists for uGMM-NN. Developing a tractable Viterbi-style procedure, typically based on a forward-backward pass, as used in Sum-Product Networks and Probabilistic Circuits [7, 9, 11], remains an open challenge for extending uGMM-NN to large-scale generative inference. However, unlike traditional graphical models, uGMM-NN avoids the need for complex Bayesian structure learning, which is generally intractable due to the combinatorial nature of graph search and the difficulty of marginal likelihood estimation. Instead, the architecture of a uGMM-NN is specified simply by choosing the number of layers and neurons, and can be trained end to end via standard backpropagation. This simplifies the model’s design and eliminates the need for expert-guided structure discovery. Hence, while generative applications remain an interesting direction for future work, the primary contribution of uGMM-NN lies in its ability to replace deterministic neurons with probabilistic units for interpretable, discriminative inference.
6 Conclusion and Future Work
This paper introduced the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel feedforward architecture that integrates probabilistic reasoning directly into the computational units of neural networks. By embedding a univariate Gaussian mixture model (uGMM) into each neuron, uGMM-NN provides a richer parameterization of activations, allowing the network to capture multimodality and quantify uncertainty in a way that standard feedforward layers cannot. This makes uGMM-NN an attractive building block for architectures that require both predictive accuracy and probabilistic interpretability.
Our preliminary experiments on the Iris and MNIST benchmarks demonstrate that uGMM-NN achieves competitive performance with standard feedforward neural networks (FFNNs) when trained discriminatively. These results highlight the feasibility of replacing conventional neurons with mixture-based alternatives, offering a new design space for deep learning architectures.
Future work will focus on several directions:
  • Extending uGMM neurons to other neural architectures such as Recurrent Neural Networks (RNNs) [16] and Transformers [17], to test their versatility across sequential and attention-based domains.
  • Developing efficient Most Probable Explanation (MPE) inference algorithms, including Viterbi-style procedures, to make uGMM-NN scalable for generative applications [11].
  • Scaling experiments to larger and multimodal datasets to evaluate robustness, generalizability, and performance under more complex real-world conditions.
In summary, uGMM-NN provides a compelling direction for embedding probabilistic modeling into deep learning architectures. With further development of inference algorithms, scalability, and theoretical analysis, uGMM-NN has the potential to bridge the gap between discriminative neural networks and generative probabilistic models, enabling architectures that are both accurate and interpretable.